A random sample of size n is drawn from a uniform population over (θ-1/3,θ+1/3),Obtain maximum likelihood estimator ofθ.


1
Expert's answer
2022-01-31T16:29:28-0500

likelihood function for a uniform distribution:

"L(\theta)=\Pi^n_{i=1}f(x_i;a,b)=\Pi^n_{i=1}\frac{1}{(b-a)^n}=\Pi^n_{i=1}\frac{1}{(b-a)^n}=\Pi^n_{i=1}\frac{1}{(\theta +1/3-\theta +1/3)^n}=\Pi^n_{i=1}\frac{1}{(2/3)^n}"


log-likelihood function:

"l(\theta)=-nlog(2/3)"


log-likelihood function does not depend on "\theta" , so MLE for "\theta" does not exist